Gio has a pack of 18 cards. Ivy has a pack of x cards. Each card has a picture of a bus or a car or a train. The number of cards with each picture in the two packs is shown in the table. Bus | Car | Train Gio's pack | 6 | 10 | 2 Ivy's pack | x-12 | 9 | 3
- (a)Write down an equation in terms of x and hence find the value of x.[4]
- (b)Find the probability that the two cards have pictures of the same type of vehicle on them.[3]
Mark scheme
Marking points from the official mark scheme (Oct/Nov 2025). The mark or mark type is shown in brackets.
- [B1] Either 6/18 * (x-12)/x Or 10/18 * 9/x seen. OE.
- [B1] Correct equation formed. Fractions may be simplified.
- [M1] Rearrange probabilities to form quadratic equation and solve to find a value for x. Condone elimination of x on denominators: 6/18 * x(x-12) = 2 * 10/18 * 9 OE. Must be an equation throughout.
- [A1] If M1 not awarded, SCB1 for [x =] 42 WWW. Note x = 42 must be selected if x = 0 is present.
- [M1] Two identified un-simplified outcomes with their x substituted. Correct values linked to identified outcomes acceptable (using x = 42).
- [M1] Add probabilities, 0 < p < 1, for 3 correct scenarios, no incorrect/repeated scenarios. Identification can be implied by un-simplified expressions.
- [A1] 0.365079... rounded to at least 3SF. OE. [SCB1] 10/42 + 5/42 + 1/126 = 23/63 OE.
Write your answer and get it marked
On Oracle Prep you can write or photograph your own answer and have it marked by AI against this mark scheme, with feedback on what you missed.
Sign up free →